OpenStudy (anonymous):

1.Solve for p 3(p+q)=p 2.solve for b then find the value of b when a=3 4a=2b-7 3. Solve for r D=rt 4. Find the width of a rectangle with a perimeter of 90 and the length of 15

OpenStudy (anonymous):

1. 3p-p=-3q therefore p=-(3/2)q 2. 2b-7=4a therefore 2b=4a+7 therefore b=2a+7/2 in order to see how much is b when a=3, just plug in a=3, and we get b=2(3)+7/2=19/2 3. r=D/t 4. perimeter=2(length)+2(width) therefore 2width=perimeter-2(length). Divide both sides by 2 and plug in numbers: width= 90/2-15=30

OpenStudy (anonymous):

Thank you 👌🏻😉

OpenStudy (anonymous):

medal?